Phantomdex:

Data collected from real estate agents were used to calculate the model ŷ = 87x + 108,620, where x represents the size of a house in square feet and ŷ represents the predicted selling price of the house in dollars. Using the model, predict the size of a house that sells for $375,000. 30,618 square feet 32,734 square feet 3,062 square feet 3,273 square feet

Phantomdex:

@jayfafr

RhianaB112:

To predict the size of a house that sells for 375,000, we can substitute 375,000 for \hat{y} in the equation \hat{y}=87x+108,620 y ^ ​ intheequation y ^ ​ =87x+108,620 and solve for x. 375,000=87x+108,620 Subtract 108,620 from both sides: 375,000-108,620=87x 266,380=87x Divide both sides by 87: x=\frac{266,380}{87} x= 87 266,380 x\approx 3,062x≈3,062

Phantomdex:

@jayfafr

RhianaB112:

what is this for ?are you trying to add someone ?

Phantomdex:

Oh no I just get a second person to check work!

RhianaB112:

@phantomdex wrote:
Oh no I just get a second person to check work!
oh ok

Luigi0210:

@rhianab112 is correct, but I'll reiterate what they said to where it's easier to read: Given \(\large y=87x+108,620 \) and y = 375000 Just plug and solve like they did: \[\Large 375000=87x+108620\] \[\Large 266380 =87x \] Divide by the 87, and you get x: \[\Large x= 3061.8 \approx 3062~ft^2\]

RhianaB112:

oh i can't believe it

RhianaB112:

i love math sm